Assume the risk free rate is 2.0%. The SP500 is considered the market. Today (TO), you invest $400 in Stock A and $600 in Stock B to create Portfolio A,B. Assume there are not taxes or dividends. The one year performance of the stocks and the market is summarized in the table below. Use this information to help answer questions Investment Total ReturnTotal Risk Beta .000 I .40 0.60 Market 12.0% 18.0% 10.0% 15.8)% 12.0% Stock B $ 600

20. In lecture we discussed how assuming the correlation between any two is zero is dangeros when assessing the total risk of a portfolio. For this problem only, assume you calculate the correlation between two assets to be 0.40 instead of 0.00. Portfolio A,Bs total risk is now closest to: 9.4% 10.2% a. b. d. 12.8% 13.2% e.
Answer #1

Weightage of A in portfolio = 0.4 and Weightage of B in portfolio = 0.6

Correlation Coefficient = 0.4

Total Risk of A = R(a) = 15 % and Total Risk of B = R(b) = 12 %

Total Risk of the Portfolio = [{Weight of A x Total Risk of A}^(2) + {Weight of B x Total Risk of B}^(2) + {2 x Weight of A x Weight of B x Total Risk of A x Total Risk of B x Correlation Coefficient}]^(1/2) = [{0.4 x 15}^(2) + {0.6 x 12}^(2) + {2 x 0.4 x 0.6 x 15 x 12 x 0.4]^(1/2) = 11.063 % ~ 11.1 %

Hence, the correct option is (c)